What type of bonds exist in each of these compounds?

Types of bonds: Ionic, nonpolar covalent, polar covalent

a) KCl

b) BCl3

c) P4

d) Br2

e) CO

f) SO2

Thanks.
Answer #1
Concepts and reason

The problem is based on the concept of the formation of the bonds can be the result of the attraction between two atoms carrying opposite charge, or it can be present due to the sharing of electron. The electron sometimes is donated to another atom, this type of bond is known as ionic bond. Sometimes, the atom cannot donate the electron and hence, they form a covalent bond.

Other than the ionic and the covalent bond some weak electrostatic bonds such as hydrogen bonds and van der Waals are also present.

A covalent bond is a bond formed between two atoms. In case, the electrostatic force between the molecules is unequally distributed. The unequal distribution of the force causes a dipole moment.

Fundamentals

A covalent bond is a bond formed between two atoms. In case, the electrostatic force between the molecules is unequally distributed. The unequal distribution of the force causes a dipole moment.

Part a

The ionic bond is a bond made by the metals such as potassium and sodium. These atoms can easily loose an electron from their outermost shell, and gain an inert configuration.

The chlorine Cl{\rm{Cl}} is an electronegative halogen. The Cl{\rm{Cl}} can easily accept an electron in its outer most shell. Together the K and Cl{\rm{Cl}} form an ionic bond, in which the metal K{\rm{K}} donates its electron to outer shell of the chlorine.

Part b

The valance shell of the boron (B)\left( {\rm{B}} \right) contains three extra electrons.

Chlorine, Cl{\rm{Cl}} tend to accept an electron and form a covalent bond. As the chlorine is highly electronegative, it tends to induce the polarization and pull the electron towards it.

Part c

The valance shell of the phosphorus (P)\left( {\rm{P}} \right) contains three extra electrons. Sometimes, it also forms covalent bond with 5{\rm{5}} atoms. In the (P4)\left( {{{\rm{P}}_{\rm{4}}}} \right) , the atoms are covalently bonded.

All the atoms belong to same element (P)\left( {\rm{P}} \right) and exhibit same electronegativity. Hence, the bonds would not exhibit any polarity.

Part d

Analysis of the valency of the bromine. The valency of the bromine can help in the determination of the number of free electrons. The bromine is a halogen and intends to accept an electron.

In the Br2{\rm{B}}{{\rm{r}}_2} , the atoms are covalently bonded. The electron present in the outermost shell of the bromine, is shared with another bromine atom to complete its shell configuration. They cannot donate electron, because of the halogens are highly electronegative.

Part e

The carbon and the oxygen exhibit three covalent bonds in the structure of CO{\rm{CO}} (carbon monoxide). The covalent bonds are the bonds, in which the electrons are neither accepted not donated.

Oxygen is an electronegative atom and pull the electrons toward it. Hence, the polarity is present between these bonds. The oxygen pulls the electron towards it. This cause the induction of the mild charge on both the atoms.

Part f

The molecule of SO2{\rm{S}}{{\rm{O}}_{\rm{2}}} (sulfur dioxide) is found to be covalently bonded.

The sulfur is less electronegative than the oxygen molecule. The sulfur and the oxygen form a bond by sharing their electrons. The oxygen, due to its high electronegative nature, pull the electron toward it and thus induces polarity.

Ans: Part a

The ionic bond is present between K{\rm{K}} and Cl{\rm{Cl}} in the compound potassium chloride (KCl)\left( {{\rm{KCl}}} \right) .